# Towards Federated Learning with on-device Training and Communication in 8-bit Floating Point

### v1

@@ C.3. Proof of the Main Theorem
where $\tau$ is a random variable that takes values in $\{1,2,\ldots,T\}$ with equal probability, $T$ is the number of rounds, $U$ is the total number of updates done in each round, the quantization scales $s_{i}$ are uniformly bounded by $S$, $w_{1}$ is the initial model, and $w_{*}$ is an optimal solution of (1).
@@
$E\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\|^{2}_{2}\leq E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right)+2S\sqrt{d}GU\eta_{t}+\eta_{t}^{2}U^{2}G^{2}+18U^{4}S\sqrt{d}G\eta_{t}^{2}L+9U^{3}\eta_{t}^{3}G^{2}L$
@@
$E\left[F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right]\leq\frac{1}{2U\eta_{t}}\left(-E\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\|^{2}_{2}+E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}\right)+S\sqrt{d}G+\frac{1}{2}\eta_{t}UG^{2}+9U^{3}S\sqrt{d}G\eta_{t}L+\frac{9}{2}U^{2}\eta_{t}^{2}G^{2}L$
@@
$E\left[\left[\frac{1}{T}\sum_{t=1}^{T}F\left(Q_{\text{rand}}\left(\mathbf{w}_{t}\right)\right)\right]-F\left(\mathbf{w}_{*}\right)\right]$ $\leq\frac{1}{2\sqrt{TU}}\left[-E\|\mathbf{w}_{T+1}-\mathbf{w}_{*}\|^{2}_{2}+E\|\mathbf{w}_{1}-\mathbf{w}_{*}\|^{2}_{2}\right]+\frac{1}{T}\frac{9}{2}UG^{2}L+\frac{1}{\sqrt{T}}G\sqrt{U}\left(\frac{1}{2}G+9U^{2}S\sqrt{d}L\right)+S\sqrt{d}G$

### v2

@@ C.3 Proof of the Main Theorem
where $\tau$ is a random variable that takes values in $\{1,2,\ldots,T\}$ with equal probability, $T$ is the number of rounds, $U$ is the total number of updates done in each round, quantization scales $s_{i}$ are uniformly bounded by $S$, $\mathbf{w}_{1}$ is the initial model, and $\mathbf{w}_{*}$ is an optimal solution of (1).
@@
$E\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\|^{2}_{2}\leq$ $E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}-2U\eta_{t}E\left(F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right)+2S\sqrt{d}GU\eta_{t}+\eta_{t}^{2}U^{2}G^{2}$
$+18U^{4}S\sqrt{d}G\eta_{t}^{2}L+9U^{3}\eta_{t}^{3}G^{2}L$
@@
$E\left[F\left(Q_{\text{rand}}\left({\mathbf{w}}_{t}\right)\right)-F\left(\mathbf{w}_{*}\right)\right]\leq$ $\frac{1}{2U\eta_{t}}\left(-E\|\mathbf{w}_{t+1}-\mathbf{w}_{*}\|^{2}_{2}+E\left\|{\mathbf{w}}_{t}-\mathbf{w}_{*}\right\|^{2}_{2}\right)$
$+S\sqrt{d}G+\frac{1}{2}\eta_{t}UG^{2}+9U^{3}S\sqrt{d}G\eta_{t}L+\frac{9}{2}U^{2}\eta_{t}^{2}G^{2}L$
@@
$E$ $\left[\left[\frac{1}{T}\sum_{t=1}^{T}F\left(Q_{\text{rand}}\left(\mathbf{w}_{t}\right)\right)\right]-F\left(\mathbf{w}_{*}\right)\right]$
$\leq\frac{1}{2\sqrt{TU}}\left[-E\|\mathbf{w}_{T+1}-\mathbf{w}_{*}\|^{2}_{2}+E\|\mathbf{w}_{1}-\mathbf{w}_{*}\|^{2}_{2}\right]+\frac{1}{T}\frac{9}{2}UG^{2}L+\frac{1}{\sqrt{T}}G\sqrt{U}\left(\frac{1}{2}G+9U^{2}S\sqrt{d}L\right)+S\sqrt{d}G$